OpenStudy (anonymous):

can someone please help me Solve x2 + 6x + 7 = 0.

pooja195 (pooja195):

Are we supposed to use the quadratic formula?

OpenStudy (anonymous):

yes @pooja195

pooja195 (pooja195):

Alright \[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] This is the formula we need to find our abc values Do you know how?

OpenStudy (anonymous):

yes b is the 6 c is the 7 but idk a @pooja195

pooja195 (pooja195):

Ok We use the standered equation to determine the abc values which is this: \[\huge~\rm~ax^2+bx+c=0\]

pooja195 (pooja195):

Match it up with your equation \[\huge~\rm~ x^2 + 6x + 7 = 0.\]

pooja195 (pooja195):

Can you tell me what the a and c values are now?

OpenStudy (anonymous):

1 and 7? @pooja195

pooja195 (pooja195):

Yes so our a=1 b=6 c=7 \[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Now we need to plug the numbers in \[\huge~\rm~x=\frac{ -(6) \pm \sqrt{(6)^2-4(1)(7)} }{ 2(1)}\]

pooja195 (pooja195):

Can you simplfy that?

OpenStudy (anonymous):

yes @pooja195

OpenStudy (anonymous):

would the -6 on the outside of the sqrt sign stay negative or posotive

pooja195 (pooja195):

it remains negative

OpenStudy (anonymous):

i didnt get what any of my answer choices were @pooja195

pooja195 (pooja195):

Thats because there is still another part

OpenStudy (anonymous):

@pooja195 ok then i think i simplified it correctly -6+_ sqrt of -16 / 2

pooja195 (pooja195):

\[\huge~\rm~x=\frac{ -6 \pm \sqrt{8} }{ 2}\]

pooja195 (pooja195):

What are your answer chocies?

OpenStudy (anonymous):

how do i take a screen shot of the window @pooja195

pooja195 (pooja195):

Im not sure... 

pooja195 (pooja195):

Is that an option?

OpenStudy (anonymous):

yes that is

pooja195 (pooja195):

Thats the simplified form of the equation i wrote above thats your final answer :)

OpenStudy (anonymous):

thank you so much

pooja195 (pooja195):

No problem :) \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)